Water from a dam falls on a turbine 30 m below at the rate of 60 m³/min. After
leaving the turbine wheel the water has a velocity of 9 m/s. What is the maximum
efficiency of the turbine? What is the maximum power developed?
1
Expert's answer
2020-11-25T05:30:21-0500

The maximum efficiency is


"\eta_\text{max}=\frac{v_\text{ptactical}}{v_\text{theoretical}}=\frac{v_\text{ptactical}}{\sqrt{gH}},\\\space\\
\eta_\text{max}=\frac{9}{\sqrt{9.8\cdot30}}=0.52"

The maximum power of the turbine is


"P_\text{max}=QgH."

Convert 60 m³/min to kg/s:


"Q=60\cdot1000/60=1000\text{ kg/s}.\\
P_\text{max}=1000\cdot9.8\cdot30=294000\text{ W}."